OpenStudy (anonymous):

Absolute Value If f(x)=-3|5x+2| and g(x)=-7 solve, a. f(x)=g(x) b.f(x)>|g(x) c. f(x)

OpenStudy (anonymous):

\[-3|5x+2|=-7\\ |5x+2|=\frac{7}{3}\\ 5x+2=\frac{7}{3}\] or \[5x+2=-\frac{7}{3}\]

OpenStudy (anonymous):

you can finish solving those two i am sure

OpenStudy (anonymous):

can you just finish the b

OpenStudy (anonymous):

is it going to be this 7/3<\x-7/3 @findingemo